OpenStudy (anonymous):

When a force of 20.0N is applied to a spring, it elongates 0.20m. Determine the period of oscillation of a 4.0kg object suspended from this spring.

OpenStudy (anonymous):

Hey there, this is my very first time doing this. Hopefully, this helps. In first place you have to consider the formula of the force of a spring: F=k(xf-x0) Therefore, 20N = K 0.20m - if you isolate K you have K= 20N/0.20m = 100N/m. You now know the constant and you have the mass of the object, so you can apply one of the formulas of the period:\[T=2\pi \sqrt{m/K}\].:: \[T=2\pi \sqrt{4kg/100N/m}\] = 0.2 s Let me know if this is helpfull for you ^^.

OpenStudy (anonymous):

the method of the above solution is correct but the answer after the values are put is 1.256